Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 22     Entry time: 09 Feb 2023, 07:14
Author: Jayanath Koliyadu 
Group: SPB 
Subject: Shift end 

X-ray delivery

  • 9.3 keV, 1-85 pulses, 1.3 mJ, stable delivery

Optical laser delivery

  • 800 nm, 1-226 pulses, stable delivery except for shift change issues

Achievements/Observations

  • PP laser alignment and optimization at IRU and PAM
  • Aligned beam until IBS
  • Checked MKB focus
  • Timing at PAM
  • First signals on GOTTHARD II

Issues

  • Issues with shift change for PP laser. Intensity flucutation observed after shift change was done. This issue was similar to what was observed last year. We will inform LAS about this issue. No LAS OCD available to call before we leave after the shift (7:30 AM).
ELOG V3.1.4-7c3fd00